In 1973 Gordon Moore, one of Intel’s founders, predicted that the number of transistors that could be placed on a single silicon chip would double every 18 months, equivalent to an annual growth of 59% (i.e., 1.59^(1.5) = 2.0). The first microprocessor was built in 1971 and had 2,250 transistors. By 2015, high-end Intel chips contained 5.5 billion transistors, nearly 2.5 million times the number of transistors 44 years earlier. What has been the annual compound rate of growth in processing power? How does it compare with the prediction of Moore’s law?
Answer #1

Answer

Formula:

A = P(1 + r)n

here A = End Value = 5.5 billion , P = initial Value = 2250 , n = number of years = 44 and we have to calculate r = compounding growth rate.

Hence we have

=> 5.5 billion = 2250(1 + r)44

=> 2.5 million = (1 + r)44

=> r = 1.39 - 1 = 0.39

Hence,  the annual compound rate of growth = 0.39.

Hence we can se see from above that actual  the annual compound rate of growth is lesser than  the prediction of Moore’s law